# Hierarchical Graph Neural Network for Compressed Speech Steganalysis

### v1

@@ 5.5 Comparison with SOTA methods
To validate the performance of the proposed model, a comparison was conducted with the following four SOTA methods: IDC [li2012detection], SS-QCCN [li2017steganalysis], RNN-SM [lin2018rnn], and CNN-LSTM [yang2019steganalysis]. These methods were selected for several reasons. First and foremost, they are specifically tailored for detecting QIM steganography, similar to our approach. Secondly, these methods are established SOTA approaches frequently used as benchmarks in various studies. Lastly, to ensure a fair and equitable comparison, these approaches were reimplemented based on their original papers, which either provided GitHub links for implementation, as with [lin2018rnn], or were thoroughly documented with sufficient details for independent reimplementation, as provided in [li2012detection, li2017steganalysis, yang2019steganalysis].
The implementations of these approaches were carried out in Python, adhering to the parameters specified in their respective papers. For IDC and SS-QCCN, given their reliance on SVM with quadratic time complexity, it’s impractical to evaluate them using the entire stego and cover segment datasets. Following the experimental settings in [li2017steganalysis], 4,000 samples were randomly selected for the training set, maintaining a cover-to-stego ratio of 1:1, and 2,000 samples for the testing set, also with a cover-to-stego ratio of 1:1. On the other hand, RNN-SM and CNN-LSTM were evaluated using the same dataset employed in our proposed approach.
@@
0.5 IDC [li2012detection] 59.15 73.35 81.90 89.40 94.05
SS-QCCN [li2017steganalysis] 61.25 77.35 87.90 93.15 95.85
RNN-SM [lin2018rnn] 70.81 84.39 93.14 96.09 97.43
CNN-LSTM [yang2019steganalysis] 71.78 85.52 92.69 96.76 98.29
@@
5 IDC [li2012detection] 68.15 84.80 95.10 98.35 99.60
SS-QCCN [li2017steganalysis] 71.30 96.15 99.40 99.90 100
RNN-SM [lin2018rnn] 67.91 88.76 96.59 98.67 99.60
CNN-LSTM [yang2019steganalysis] 89.98 96.80 99.66 99.81 99.87
@@
10 IDC [li2012detection] 64.50 85.60 95.95 99.55 99.70
SS-QCCN [li2017steganalysis] 72.15 98.75 99.80 100 100
RNN-SM [lin2018rnn] 62.41 92.74 98.32 99.65 99.93
CNN-LSTM [yang2019steganalysis] 92.38 99.65 99.74 99.93 99.96
@@
To further validate the approach, a detection comparative analysis was conducted with four recently proposed methods: E-SWAN [wang2025swan], Bi-LSTM-3DCNN [li2025compressed], TENet [zhang2023tenet], and FedSpey [tian2023fedspy]. These methods use the same dataset and QIM steganography technique as our work. Due to the complexity of the techniques and limited available information, these methods were not re-implemented. Consequently, the comparison is based on results presented in the corresponding published papers and is limited to scenarios common to our study.
@@ 5.6 Time efficiency and compuational complexity
Table 7 provides insights into the time efficiency of our approach compared to SOTA methods in three sample length scenarios (0.5 seconds, 5 seconds, and 10 seconds). The detection time is reported as the mean value [$\pm$ standard deviation]. Our approach consistently demonstrates superior time efficiency, outperforming the SOTA methods in terms of speed across all sample lengths. The IDC approach presents closer DTs to our approach, which can be explained by the simple architecture adopted in IDC, focusing only on inter-frame correlation. SS-QCCN, on the other hand, yields the least favourable results, primarily due to its method of extracting intra- and inter-frame correlations and its utilization of the PCA method to reduce the dimensionality of a 131,072-dimensional feature vector to 300. This process incurs a significant amount of time, contributing to the observed extended DT in this method. For CNN-LSTM and RNN-SM, the complicated architecture, particularly for CNN-LSTM, which involves sequential and convolutional operations, might be more computationally intensive, especially when dealing with longer sequences, leading to slower DTs compared to ours. These results demonstrate the exceptional efficiency of our proposed model, enabling near real-time identification of hidden data in VoIP voice transmissions. This efficiency has significant practical implications, allowing for continuous monitoring of communications and immediate detection of steganographic content. The model’s scalability across various sample lengths and its superior speed compared to SOTA methods indicate excellent adaptability and resource efficiency. This performance could significantly enhance operational efficiency in security applications and allow for potential seamless integration into existing VoIP infrastructure.
Table 7: DT (s) of our approach compared to SOTA method. Results are reported as Mean [$\pm$ standard deviation]
@@
IDC [li2012detection] 0.019 [$\pm$ 0.004] 0.037 [$\pm$ 0.003] 0.049 [$\pm$ 0.002]
SS-QCCN [li2017steganalysis] 0.099 [$\pm$ 0.056] 0.134 [$\pm$ 0.070] 0.146 [$\pm$ 0.053]
RNN-SM [lin2018rnn] 0.032 [$\pm$ 0.057] 0.081 [$\pm$ 0.083] 0.127 [$\pm$ 0.075]
CNN-LSTM [yang2019steganalysis] 0.075 [$\pm$ 0.069] 0.119 [$\pm$ 0.068] 0.152 [$\pm$ 0.083]
Our 0.016 [$\pm$ 0.001] 0.029 [$\pm$ 0.022] 0.045 [$\pm$ 0.035]
While our reported DT in Table 7 and Figure 12 focuses on single-sample inference—a common practice in related steganalysis works [yang2019hierarchical, yang2022real, li2022general, lin2018rnn, yang2019steganalysis, li2012detection, li2017steganalysis, wu2020steganalysis, yang2018steganalysis, yang2019fast, hu2021detection, wang2021fast, li2021detection, wang2022steganalysis, li2023sanet, wang2025swan, li2025compressed, yang2020fcem, qiu2022steganalysis, zhang2023tenet, tian2023fedspy, zhang2024spm] that assesses minimal latency for individual streams—the architecture of our proposed GraphSAGE-based model is inherently well-suited for batch processing, a critical aspect for high-throughput real-world deployments. The core operations within our GraphSAGE layers are highly parallelizable, meaning multiple sliding windows (each representing a speech segment) can be grouped into batches and processed simultaneously. This capability, efficiently managed by modern deep learning frameworks on GPUs or multi-core CPUs, allows for a significant reduction in the effective detection time per sample when handling a high volume of traffic. Thus, the low single-sample DT provides a strong foundation for our system’s scalability and high throughput in large-scale, real-time VoIP environments.
@@
IDC [li2012detection] 735K 1.53M
SS-QCCN [li2017steganalysis] 5.2M 11.63M
RNN-SM [lin2018rnn] 33K 6.1M
CNN-LSTM [yang2019steganalysis] 258K 46.2M
@@ 6 Real-world applications, limitations, and improvements
Although our approach demonstrates strong performance, it faces some limitations and challenges. Firstly, the model struggles with very short sample lengths (less than 0.5 seconds) and extremely low embedding rates (below 20%). This difficulty arises because there is limited information available in these cases, making it challenging to extract enough features for accurate detection. One possible solution to overcome this limitation is to integrate GAT [velickovic2017graph] into the model. GATs are a type of GNN that utilizes attention mechanisms to focus selectively on the most relevant connections and information within the graph. By assigning higher weights to connections that reveal significant variations in QIM sequences, GATs ”could be” more effectively capture subtle changes in short samples and low embedding rates.

### v2

@@ 5.5 Comparison with SOTA methods
To validate the performance of the proposed model, a comparison was conducted with the following four SOTA methods: IDC [36], SS-QCCN [37], RNN-SM [16], and CNN-LSTM [34]. These methods were selected for several reasons. First and foremost, they are specifically tailored for detecting QIM steganography, similar to our approach. Secondly, these methods are established SOTA approaches frequently used as benchmarks in various studies. Lastly, to ensure a fair and equitable comparison, these approaches were reimplemented based on their original papers, which either provided GitHub links for implementation, as with [16], or were thoroughly documented with sufficient details for independent reimplementation, as provided in [36, 37, 34].
The implementations of these approaches were carried out in Python, adhering to the parameters specified in their respective papers. For IDC and SS-QCCN, given their reliance on SVM with quadratic time complexity, it’s impractical to evaluate them using the entire stego and cover segment datasets. Following the experimental settings in [37], 4,000 samples were randomly selected for the training set, maintaining a cover-to-stego ratio of 1:1, and 2,000 samples for the testing set, also with a cover-to-stego ratio of 1:1. On the other hand, RNN-SM and CNN-LSTM were evaluated using the same dataset employed in our proposed approach.
@@
0.5 IDC [36] 59.15 73.35 81.90 89.40 94.05
SS-QCCN [37] 61.25 77.35 87.90 93.15 95.85
RNN-SM [16] 70.81 84.39 93.14 96.09 97.43
CNN-LSTM [34] 71.78 85.52 92.69 96.76 98.29
@@
5 IDC [36] 68.15 84.80 95.10 98.35 99.60
SS-QCCN [37] 71.30 96.15 99.40 99.90 100
RNN-SM [16] 67.91 88.76 96.59 98.67 99.60
CNN-LSTM [34] 89.98 96.80 99.66 99.81 99.87
@@
10 IDC [36] 64.50 85.60 95.95 99.55 99.70
SS-QCCN [37] 72.15 98.75 99.80 100 100
RNN-SM [16] 62.41 92.74 98.32 99.65 99.93
CNN-LSTM [34] 92.38 99.65 99.74 99.93 99.96
@@
To further validate the approach, a detection comparative analysis was conducted with four recently proposed methods: E-SWAN [46], Bi-LSTM-3DCNN [47], TENet [49], and FedSpey [52]. These methods use the same dataset and QIM steganography technique as our work. Due to the complexity of the techniques and limited available information, these methods were not re-implemented. Consequently, the comparison is based on results presented in the corresponding published papers and is limited to scenarios common to our study.
@@ 5.6 Time efficiency and compuational complexity
Table 8 provides insights into the time efficiency of our approach compared to SOTA methods in three sample length scenarios (0.5 seconds, 5 seconds, and 10 seconds). The detection time is reported as the mean value $\pm$ standard deviation]. Our approach consistently demonstrates superior time efficiency, outperforming the SOTA methods in terms of speed across all sample lengths. The IDC approach presents closer DTs to our approach, which can be explained by the simple architecture adopted in IDC, focusing only on inter-frame correlation. SS-QCCN, on the other hand, yields the least favourable results, primarily due to its method of extracting intra- and inter-frame correlations and its utilization of the PCA method to reduce the dimensionality of a 131,072-dimensional feature vector to 300. This process incurs a significant amount of time, contributing to the observed extended DT in this method. For CNN-LSTM and RNN-SM, the complicated architecture, particularly for CNN-LSTM, which involves sequential and convolutional operations, might be more computationally intensive, especially when dealing with longer sequences, leading to slower DTs compared to ours. These results demonstrate the exceptional efficiency of our proposed model, enabling near real-time identification of hidden data in VoIP voice transmissions. This efficiency has significant practical implications, allowing for continuous monitoring of communications and immediate detection of steganographic content. The model’s scalability across various sample lengths and its superior speed compared to SOTA methods indicate excellent adaptability and resource efficiency. This performance could significantly enhance operational efficiency in security applications and allow for potential seamless integration into existing VoIP infrastructure.
Table 8: DT (s) of our approach compared to SOTA method. Results are reported as Mean $\pm$ standard deviation]
@@
IDC [36] 0.019 $\pm$ 0.004] 0.037 $\pm$ 0.003] 0.049 $\pm$ 0.002]
SS-QCCN [37] 0.099 $\pm$ 0.056] 0.134 $\pm$ 0.070] 0.146 $\pm$ 0.053]
RNN-SM [16] 0.032 $\pm$ 0.057] 0.081 $\pm$ 0.083] 0.127 $\pm$ 0.075]
CNN-LSTM [34] 0.075 $\pm$ 0.069] 0.119 $\pm$ 0.068] 0.152 $\pm$ 0.083]
Our 0.016 $\pm$ 0.001] 0.029 $\pm$ 0.022] 0.045 $\pm$ 0.035]
While our reported DT in Table 8 and Figure 12 focuses on single-sample inference—a common practice in related steganalysis works [13, 14, 15, 16, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 18, 49, 52, 54] that assesses minimal latency for individual streams—the architecture of our proposed GraphSAGE-based model is inherently well-suited for batch processing, a critical aspect for high-throughput real-world deployments. The core operations within our GraphSAGE layers are highly parallelizable, meaning multiple sliding windows (each representing a speech segment) can be grouped into batches and processed simultaneously. This capability, efficiently managed by modern deep learning frameworks on GPUs or multi-core CPUs, allows for a significant reduction in the effective detection time per sample when handling a high volume of traffic. Thus, the low single-sample DT provides a strong foundation for our system’s scalability and high throughput in large-scale, real-time VoIP environments.
@@
IDC [36] 735K 1.53M
SS-QCCN [37] 5.2M 11.63M
RNN-SM [16] 33K 6.1M
CNN-LSTM [34] 258K 46.2M
@@ 6 Real-world applications, limitations, and improvements
Although our approach demonstrates strong performance, it faces some limitations and challenges. Firstly, the model struggles with very short sample lengths (less than 0.5 seconds) and extremely low embedding rates (below 20%). This difficulty arises because there is limited information available in these cases, making it challenging to extract enough features for accurate detection. One possible solution to overcome this limitation is to integrate GAT [63] into the model. GATs are a type of GNN that utilizes attention mechanisms to focus selectively on the most relevant connections and information within the graph. By assigning higher weights to connections that reveal significant variations in QIM sequences, GATs ”could be” more effectively capture subtle changes in short samples and low embedding rates.